OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of f on the given interval. (Round all answers to two decimal places.) f(x) = x - ln(8x) [0.5, 2] Find the absolute maximum and absolute minimum values of f on the given interval. (Round all answers to two decimal places.) f(x) = x - ln(8x) [0.5, 2] @Mathematics

OpenStudy (anonymous):

Erm, so I suppose i start with \[f'(x)=1-1/x\] And...now i'm stumped...

OpenStudy (anonymous):

take the derivative of the function. set it equat to zero and solve for x. this gives local max and min x values. to find the y's plug those x s back into ito the original function

OpenStudy (anonymous):

don't forget to plug the 0.5 and 2 into the original functiontoo. the smallest f(x) will be the absmin and the biggest will be abs max

OpenStudy (anonymous):

Okay so f'x=1-1/1 =0 x=1 So I plug that back into original function? f(x)=1-ln(8)????

OpenStudy (anonymous):

Wait, why would i plug in .5 and 2. The abs min or max could be 1.232 for all I know?

OpenStudy (anonymous):

yes. also the endpoints of the interval. of the all the f(x) the largest would be the max and the smallest would be the min

OpenStudy (anonymous):

Yes to what? I don't see what you are trying to say. Okay so I plugged x=1 into the original function like you said. f(x)=1-ln(8). What in the world do i do now? I don't know what ln(8) is without a calculator...

OpenStudy (anonymous):

ln8 = 2.07

OpenStudy (anonymous):

Did you figure that out in your head...? I'm supposed to be able to do all of this without a calculator.

OpenStudy (anonymous):

But regardless, 1-2.07=-1.07 Okay, that was correct! Wow... So how do I find the max now?

OpenStudy (anonymous):

but the ln function increases so you know ln 4< ln8<16

OpenStudy (anonymous):

the largest solution is the max

OpenStudy (anonymous):

Got it, -.77 Okay so basically I just have to actually picture what the function looks like in my head. Thanks!

OpenStudy (anonymous):

yes you must include the endpoints because they can be actually the min or max and the derivative only gives you a relative max or min